OpenStudy (anonymous):

After a dilation with a center of (0, 0), a point was mapped as (2, –5) → (3, y). A student determined y to be –2.5. Evaluate the student's answer. A. The student is correct. B. The student incorrectly calculated the scale factor as 0.5. C. The student incorrectly divided by the scale factor instead of multiplying by it. D. The student incorrectly added the scale factor instead of multiplying by it.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

a = factor of dilation 3 = 2*a and y = -5*a From the first coordinate, a = 3/2, therefore y = -5*3/2 = -7.5 not -2.5 If the factor "a" were 0.5 then y would be -5*0.5 = -2.5(the result he got) therefore B. should be the correct answer